PartnerObserving Log for 2007 June 11/12

Observers: V. Burwitz, D. Thompson

Telescope Operator: D. Gonzalez

Instrument: LBC-Blue

[UT]

00:33 so far bad weather thick fog light snow on dome

11:07 Dome Remained Closed even though sky cleared as melting snow kept water continuously flowing down sides of Dome.

11:30 in twightlight many high clouds above observatory.

-- VadimBurwitz - 13 Jun 2007
Topic revision: r5 - 21 Jul 2009, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback